Find an integrating factor of the form Xy and solve the equation. (2x-172-9y)dx + (3y-6x) dy=0, y(1) =1 OA 4x2y3 – 3x3y2 =
Answer #1

) ymt - 60 scht Amo (2x-92-94) dx + (3 7-63) dy 91-1 • Multiply the equahon by M (any) to get (uid,y) = xhym) = (2x y² Xx^yand xh • Now matching conreuponding terms jie, analy xhym -) 2 (m+2) = 3n and -9(m + 1) = -6 (nt) - 2m +4-35 -0 qmtq - 60-6=(A 3 2 So, the integranng fachor isbry Now put values of n and min (2 x y 3 - 9x²42) dx + (3x² y ² - 6x²y) dy = 0 Solution to